# Effects of orbital eccentricity on continuous gravitational waveforms from triaxially-deformed precessing neutron stars in tight binaries

### v1

@@ IV.1 Signal model in detector frame
where $F_{+,\times}(t)$ are the antenna pattern functions of the GW detector (see Ref. Jaranowski et al. (1998) for the explicit expressions). The functions $F_{+,\times}(t)$ depend on the following parameters: $\gamma_{\rm o}$, which defines the detector’s orientation to the local geographical directions; $\zeta$, the angle between the arms of the interferometer; $\lambda$, the geographical latitude of the detector location; $(\alpha,\delta)$, the right ascension and declination of the source; $\psi_{\rm p}$, the GW polarization angle; $\Omega_{\rm Er}$, the Earth’s rotational angular frequency; and $\phi_{\rm r}$, the initial phase of the Earth’s diurnal rotation.
@@ IV.2 Effects of eccentricity on parameter estimation
The FIM is numerically calculated using Mathematica with parameters in Table 1 and Table 2, for different binary parameters $(P_{\rm b},e)$. For the COMPAS model, the orbital eccentricity is set to $e\in\big{\{}0.01,0.03,0.05,0.1,0.18,0.30\big{\}}$, and the most likely values of the orbital period distribution are $P_{\rm b}\in\big{\{}1535,1535,1535,1838,2200,2634\big{\}}~{}{\rm s}$. In terms of the IsolatedFast and DynamicalFast fast-merging DNS evolutionary channels, the eccentricities are set to $e\in\big{\{}0.1,0.15,0.2,0.25,0.3\big{\}}$ and $e\in\big{\{}0.6,0.65,0.7,0.75,0.8\big{\}}$, respectively, and the corresponding orbital period for both scenarios is assumed to be the mean value of the distribution, which is $5155~{}{\rm s}$ (see Fig. 3).
Figure 4: The SNR as a function of eccentricity is estimated for the first three waveform components, $h_{1}(t),h_{2}(t),h_{3}(t)$, from a spinning NS in different DNS models using the Cosmic Explorer. The parameters used here can be found in Table 1 and Table 2.
@@
The SNR as a function of eccentricity for the first three waveform components from a spinning NS in different binaries is shown in Fig. 4, where, from left to right, subplots correspond to the waveform components $h_{1}(t)$, $h_{2}(t)$, and $h_{3}(t)$, respectively. In the figure, the blue, orange, and magenta colors represent the DNS population from COMPAS model, IsolatedFast model, and DynamicalFast model, respectively. The three colors are the same as in Fig. 3 and Fig. 5 below, for ease of comparison. For the waveform component $h_{1}(t)$, the SNR of the NS waveforms in DNSs from the COMPAS model decreases as eccentricity increases, whereas in the IsolatedFast and DynamicalFast models, the SNR increases with increasing eccentricity. The results for the waveform components $h_{2}(t)$ and $h_{3}(t)$ are different.
@@ V Conclusions
In this study, we calculate the continuous GW waveforms emitted by spinning NSs in binaries that experience spin precession due to spin-orbit coupling, incorporating the influence of orbital eccentricity. We compared these waveforms with those emitted by NSs in circular binary systems. By calculating the time required for the FF to decrease to 0.97 for the first three NS waveform components in binaries, we found that for DNS systems generated by the COMPAS and IsolatedFast models, where the orbital eccentricities are relatively small, there is no significant impact of orbital eccentricity on the waveforms emitted by spinning NSs in binaries, indicating the waveforms without eccentricity is sufficiently accurate in the detection of these sources. However, for DNS systems generated by the DynamicalFast model, where the orbital eccentricities are larger, the orbital eccentricity has a more significant impact on the waveforms. Additionally, by accounting for the Doppler modulation effect introduced by eccentric orbital motion, we assessed the SNR and the accuracy of parameter estimation for NS waveforms across various DNS models using the Cosmic Explorer detector as an example. For NS waveforms within the COMPAS-based DNS population, even though the SNR detected by the Cosmic Explorer decreases with increasing eccentricity, the orbital period of this population increases with eccentricity. Consequently, the parameter estimation for $\ln P_{\rm b}$, $\ln e$, $\cos\iota$, and $\theta_{S}$ improves with increasing eccentricity. There remains an open question on the number of detectable NSs emitting continuous GWs in eccentric orbits, which we leave for future studies.
@@ Acknowledgements.
Y.W. gratefully acknowledges support from the National Key Research and Development Program of China (No. 2023YFC2206702 and No. 2022YFC2205201), the National Natural Science Foundation of China (NSFC) under Grants No. 11973024, Major Science and Technology Program of Xinjiang Uygur Autonomous Region (No. 2022A03013-4), and Guangdong Major Project of Basic and Applied Basic Research (Grant No. 2019B030302001). L.S. acknowledges support from the Beijing Natural Science Foundation (1242018), the National SKA Program of China (2020SKA0120300), the National Natural Science Foundation of China (11991053), the Max Planck Partner Group Program funded by the Max Planck Society and the High-performance Computing Platform of Peking University.

### v2

@@ IV.1 Signal model in detector frame
where $F_{+,\times}(t)$ are the antenna pattern functions of the GW detector (see Ref. Jaranowski et al. (1998) for the explicit expressions). The functions $F_{+,\times}(t)$ depend on the following parameters: $\gamma_{\rm o}$, which defines the detector’s orientation to the local geographical directions; $\zeta$, the angle between the arms of the interferometer; $\lambda$, the geographical latitude of the detector location; $(\alpha,\delta)$, the right ascension and declination of the source; $\psi_{\rm p}$, the GW polarization angle; $\Omega_{\rm Er}$, Earth’s rotational angular frequency; and $\phi_{\rm r}$, the initial phase of Earth’s diurnal rotation.
@@ IV.2 Effects of eccentricity on parameter estimation
Figure 4: The SNR as a function of eccentricity is estimated for the first three waveform components, $h_{1}(t),h_{2}(t),h_{3}(t)$, from a spinning NS in different DNS models using the Cosmic Explorer. The parameters used here can be found in Tables 1 and 2.
@@
The FIM is numerically calculated using Mathematica with parameters in Tables 1 and 2, for different binary parameters $(P_{\rm b},e)$. For the COMPAS model, the orbital eccentricity is set to $e\in\big{\{}0.01,0.03,0.05,0.1,0.18,0.30\big{\}}$, and the most likely values of the orbital period distribution are $P_{\rm b}\in\big{\{}1535,1535,1535,1838,2200,2634\big{\}}~{}{\rm s}$. In terms of the IsolatedFast and DynamicalFast fast-merging DNS evolutionary channels, the eccentricities are set to $e\in\big{\{}0.1,0.15,0.2,0.25,0.3\big{\}}$ and $e\in\big{\{}0.6,0.65,0.7,0.75,0.8\big{\}}$, respectively, and the corresponding orbital period for both scenarios is assumed to be the mean value of the distribution, which is $5155~{}{\rm s}$ (see Fig. 3).
@@
The SNR as a function of eccentricity for the first three waveform components from a spinning NS in different binaries is shown in Fig. 4, where, from left to right, subplots correspond to the waveform components $h_{1}(t)$, $h_{2}(t)$, and $h_{3}(t)$, respectively. In the figure, the blue, orange, and magenta colors represent the DNS population from the COMPAS, IsolatedFast, and DynamicalFast models, respectively. The three colors are the same as in Figs. 3 and 5 below, for ease of comparison. For the waveform component $h_{1}(t)$, the SNR of the NS waveforms in DNSs from the COMPAS model decreases as eccentricity increases, whereas in the IsolatedFast and DynamicalFast models, the SNR increases with increasing eccentricity. The results for the waveform components $h_{2}(t)$ and $h_{3}(t)$ are different.
@@ V Conclusions
In this study, we calculate the continuous GW waveforms emitted by spinning NSs in binaries that experience spin precession due to spin-orbit coupling, incorporating the influence of orbital eccentricity. We compared these waveforms with those emitted by NSs in circular binary systems. By calculating the time required for the FF to decrease to 0.97 for the first three NS waveform components in binaries, we found that for DNS systems generated by the COMPAS and IsolatedFast models, where the orbital eccentricities are relatively small, there is no significant impact of orbital eccentricity on the waveforms emitted by spinning NSs in binaries, indicating the waveforms without eccentricity is sufficiently accurate in the detection of these sources. However, for DNS systems generated by the DynamicalFast model, where the orbital eccentricities are larger, the orbital eccentricity has a more significant impact on the waveforms. Additionally, by accounting for the Doppler modulation effect introduced by eccentric orbital motion, we assessed the SNR and the accuracy of parameter estimation for NS waveforms across various DNS models using the Cosmic Explorer detector as an example. For NS waveforms within the COMPAS-based DNS population, even though the SNR detected by the Cosmic Explorer decreases with increasing eccentricity, the orbital period of this population increases with eccentricity. Consequently, the parameter estimation for $\ln P_{\rm b}$, $\ln e$, $\cos\iota$, and $\theta_{S}$ improves with increasing eccentricity. There remains an open question on the number of detectable NSs emitting continuous GWs in eccentric orbits, which we leave for future studies. We note that many current continuous wave searches use lower fitting factors due to computational limitations Mukherjee et al. (2023). Also, torque effects limit coherent integration times to several days for accreting binary systems like Sco X-1 Zhang et al. (2021). These practical considerations should be taken into account in future work.
@@ Acknowledgements.
We thank the anonymous referee for helpful comments and suggestions. W.-F.F. is supported by the National Natural Science Foundation of China under Grant No. 12447109. T.L. is supported by the China Postdoctoral Science Foundation Grant No. 2024M760692. Y.W. gratefully acknowledges support from the National Key Research and Development Program of China (No. 2023YFC2206702 and No. 2022YFC2205201), the National Natural Science Foundation of China under Grant No. 11973024, Major Science and Technology Program of Xinjiang Uygur Autonomous Region (No. 2022A03013-4), and Guangdong Major Project of Basic and Applied Basic Research (Grant No. 2019B030302001). L.S. acknowledges support from the Beijing Natural Science Foundation (1242018), the National SKA Program of China (2020SKA0120300), the National Natural Science Foundation of China (11991053), the Max Planck Partner Group Program funded by the Max Planck Society, and the High-Performance Computing Platform of Peking University.